what is the highest order spectrum which may be seen with monochromatic light of wavelength 6000A by means of a grating spectrum 5000per lines



1
Expert's answer
2021-07-14T13:41:18-0400
  • Diffraction can be given by the formula: nλ=dsinθ
  • To get M=Mmax,  θ has to be  θmax=90o
  • Since we have 1/d as 5000, we can replace these in an equation and solve
  • Mmax=d/λ=1/5000x1/6x10-5. This gives Mmax=3.33 hence Mmax can be rounded off to the nearest whole number as Mmax=3.